A solution containing 12.0% sodium hydroxide by mass has a density of 1.131g/ml. What volume of this solution contain 5.00 mol of NaOH

Grade:11

1 Answers

Arun
4 years ago

Let the required volume be ‘V’ ml.

Mass of V ml of the solution = density*volume = 1.131V g.

Mass of NaOH = (12/100)*1.131V = 0.136V g.

Moles of NaOH = given mass/molar mass = 0.136V/40 = 0.0034V.

Given- the required volume contains 5 mol NaOH.

0.0034V = 5

V =1472.93 ml